# Devising PoPStat: A metric bridging population pyramids with global disease mortality

### v1

@@ 3 Results
3.1 Comparison of PoPStat with other indicators
In Figure 6, ischaemic heart disease (IHD) has been used as an example to calculate the correlations of other well-recognized demographic and socio-economic indicators with mortality rate and to compare their performance with PoPStat. As seen in Figure 6(a), the linear variation of the IHD mortality rate has been captured best by the PoPStat with a moderate correlation of - 0.74 (Latvia, p: < 0.001). Meanwhile, the median age also had a moderate correlation of 0.70 (p: <0.001) with mortality due to IHD but it was lower than the PoPStat. As illustrated in Figure 3, PoPStat also showed a better correlation with IHD mortality than the correlations with life expectancy, population density, HDI, and GDP per capita.
@@ 3.1 Comparison of PoPStat with other indicators
3.2 PoPStat and disease mortality
3.2.1 Level 1 Disease Categories
Among the three aggregated disease categories in Level 1, the highest PoPStat was observed for non-communicable diseases (NCD) (PoPStat: - 0.84, reference country: Japan, p-value: <0.001) indicating a statistically significant, strong correlation with the underlying shape of the population pyramid across countries. As shown in Figure 7(a), countries with constrictive and more symmetrically shaped population pyramids such as Japan, Italy, Portugal and Greece had higher mortality due to NCDs. The PoPDivergence – which, in this situation, represents the divergence of population from a constrictive pyramid towards an expansive pyramid - explained $71.6\%$ of the variance in global mortality attributed to NCDs.
The lowest PoPStat was observed for injuries (0.29, Singapore, p: <0.001) suggestive of a weak correlation with the demographic structure of the country. Countries like Singapore, Mauritius, Chile and Saint Lucia with constructive pyramids that were largely symmetrical had lower injury-related mortality. However, the deviation from these pyramids towards more progressive shapes explained only 8.4% of the variation in injury-related mortality. As seen in Figure 7(b), the aggregate of communicable, maternal, neonatal and nutritional diseases (CMNN) had a PoPStat of 0.50 (Singapore, p <0.001), indicating a moderate association with the shape of the population pyramid. Countries with more symmetrical and constrictive pyramids like Singapore, Mauritius, Chile and Saint Lucia saw lower CMNN death rates. The divergence from these regressive population structures toward progressive pyramids accounted for nearly one-fourth of the change in mortality attributed to CMNN causes.
(a) Non-Communicable Diseases
@@ 3.2.1 Level 1 Disease Categories
(b) Communicable, maternal, neonatal, and nutritional diseases
(c) Cardiovascular Diseases
(d) Neglected Tropical Diseases
Figure 7: Main caption describing the entire figure
3.2.2 Level 2 Disease Categories
Non-Communicable Diseases
Among NCDs, the largest PoPStat (0.90, p: <0.001) was noted for neurological diseases, with Zambia as the reference population pyramid. Similar countries like Burkina Faso, Cameroon and Yemen - with expansive, predominantly symmetrical population pyramids - had a lower burden of deaths due to neurological diseases. Shifting away from these early stages of the demographic transition explained 81.4% of the variation in mortality assigned to neurological causes. The next highest PoPStat was reported for neoplasms (-0.90, Japan, p: <0.001). Populations with symmetrical, constrictive pyramids - for example, Japan, Italy, Portugal and Greece - were associated with higher cancer mortality rates in the world. 80.4% of the change in cancer mortality could be explained by the PoPDivergence from a constrictive reference pyramid. Cardiovascular diseases (CVD) had a strong PoPStat of – 0.73 (p: <0.001), with Japan as the reference population. Figure 7(c) illustrates how countries with symmetrical and constrictive population pyramids identified before had more death rates due to cardiovascular diseases as well. The variance in cardiovascular mortality explained by the divergence of populations from constrictive pyramids towards expansive pyramids was 53.4%. Meanwhile, mental disorders also had a strong PoPStat of – 0.70 (Japan, p: <0.001). Similar countries with symmetrically constrictive population pyramids shared a higher burden and shifting from them toward more progressive shapes explained nearly half of the change in deaths due to mental illnesses
Musculoskeletal disorders had a moderate PoPStat of – 0.64 (Japan, p: <0.001). Deaths due to musculoskeletal disorders were also associated with similar constrictive population pyramids and divergence from them explained 41.8% of the change in mortality. The PoPStat for chronic respiratory diseases was -0.53 (p<0.001) with France as the reference country. Constrictive population pyramids like Finland, Belgium and the UK which were less inverted than former examples saw higher death rates due to chronic respiratory diseases. Deviation from this demographic structure explained 28.1 of the change in fatalities due to chronic respiratory illnesses. Substance use disorders also showed a moderate PoPStat of -0.53 (p: < 0.001) with Japan as the reference. Countries like Italy, Portugal and Greece, characterized by symmetrical constrictive pyramids, experienced higher mortality. The divergence from these pyramids explained 28.3% of the variance in substance-related deaths. Digestive diseases had a moderate PoPStat of 0.50 (p: <0.001) with Qatar as the reference population pyramid. Countries with constrictive pyramids that were asymmetrical or male-predominated – like the United Arab Emirates, Oman and Maldives – had lower mortality due to digestive diseases. PoPDivergence in relation to such pyramids explained nearly one-fourth of the variation in deaths caused by digestive disorders. The mortality due to other NCDs category (0.49, Singapore, p: <0.001) was associated with symmetrical and constrictive population pyramids. PoPDivergence from these pyramids accounted for 23.4% of the change in deaths due to other NCDs. Meanwhile, diabetes and kidney diseases maintained the lowest PoPStat of -0.41 (p: <0.001), with Argentina becoming the reference population. Similar countries with late-expansive population pyramids – Costa Rica, Sri Lanka and St. Vincent and Grenadines – had higher mortality attributed to diabetic and kidney diseases. Nonetheless, the change in diabetic deaths explained by the departures from a late-expansive pyramid was 16.6%.
@@ 3.2.2 Level 2 Disease Categories
As visualized in Figure 7(d), the highest PoPStat among communicable diseases was observed for neglected tropical diseases (NTD) (0.83, Malta, p: < 0.001), indicating a very strong association between these diseases and the demographic structure. Countries with symmetrical, constrictive population pyramids such as Malta, Poland, Cyprus and Slovakia experienced lower mortality from NTDs. The PoPDivergence—reflecting movement from constrictive pyramids toward expansive pyramids—explained 69.7% of the variance in mortality related to these conditions. A similar pattern emerged for other infectious diseases which had a very strong PoPStat of 0.82 (p: < 0.001). Malta again served as the reference population with countries like Poland, Cyprus and Slovakia exhibiting lower mortality rates due to other infections. The shift from these symmetrically constrictive pyramids towards expansive shapes amounted to 68.4% of the variation in mortality from other infectious diseases.
Enteric infections demonstrated a strong PoPStat of 0.73 (p: < 0.001). Similar countries like Malta, characterized by symmetrical and constrictive pyramids, had lower mortality rates. The divergence from these structures towards early-stage pyramids explained 54.0% of the variance in enteric infection-related mortality. HIV and sexually transmitted infections (STIs) showed a moderate PoPStat of -0.64 (p: < 0.001) with the Central African Republic as the reference country. Countries with symmetrical and expansive demographic pyramids such as Niger, Somalia and Chad exhibited higher mortality rates. Deviations away from these early-stage pyramids explained 41.5% of the change in mortality. Respiratory infections and tuberculosis had the lowest PoPStat among communicable diseases (0.14, Qatar, p: < 0.001), suggesting a weak association with demographic structures. Asymmetrically constrictive pyramids like Qatar, UAE, Bahrain and Maldives had a lower burden of deaths. However, divergence from these pyramids only accounted for 2.0% of the variance in mortality.
Maternal, neonatal and nutritional diseases
Maternal and neonatal conditions exhibited a very strong PoPStat 0f -0.91 (p < 0.001) with the Central African Republic as the reference. Countries such as Niger, Somalia and Chad which shared expansive population pyramids saw higher death rates. Deviations from these structures explained 83.6% of the variance in maternal and neonatal mortality. Meanwhile, nutritional deficiencies showed a moderate PoPStat of 0.56 (Cyprus, p: < 0.001). Countries like Luxembourg, Chile and Iceland - having symmetrical, constrictive pyramids- exhibited lower mortality rates. Transitioning away from these pyramids towards expansive ones explained 32.3% of the change in mortality.
@@
Transport injuries (TI) (-0.62, p: < 0.001) indicated a moderate association with the underlying demography across countries. Burundi served as the reference population with countries like Mali, Burkina Faso and Mozambique showing similar symmetrically expansive pyramids and higher mortality. Movement away from these population pyramids accounted for 38.5% of the variance in mortality. Unintentional injuries (UI) displayed a moderate PoPStat of 0.52 (p: < 0.001). Kuwait was the reference population and countries with both expansive and stationary pyramids like Brunei, Iran and Bahrain showed higher mortality rates. The PoPDivergence explained 27.4% of the variation. Self-harm and interpersonal violence had a very weak PoPStat of 0.18 (p: < 0.001). Qatar was the reference and similar asymmetrically constrictive pyramids like Bahrain and the UAE had lower mortality rates. The shift in demographic structures accounted for only 3.2% of the variation of mortality due to self-harm and interpersonal violence.
@@
Disease Reference PoPStat $r^{2}$ p-value Closest population 2nd Closest population 3rd Closest population
Ischemic heart disease Latvia -0.751 0.562 <0.001 Estonia Lithuania Slovenia
Stroke France -0.497 0.247 <0.001 Belgium United Kingdom Finland
Pulmonary Arterial Hypertension Oman 0.375 0.14 <0.001 Saudi Arabia Maldives Bahrain
Chronic obstructive pulmonary disease France -0.621 0.386 <0.001 Belgium United Kingdom Finland
Asthma Central African Republic -0.594 0.352 <0.001 Niger Somalia Chad
Breast cancer Japan -0.815 0.664 <0.001 Italy Portugal Greece
Colon and rectum cancer Japan -0.922 0.85 <0.001 Italy Portugal Greece
Cervical cancer Tonga -0.257 0.066 <0.001 Samoa Syria Comoros
Prostate cancer Japan -0.716 0.513 <0.001 Italy Portugal Greece
Liver cancer Oman 0.499 0.249 <0.001 Saudi Arabia Maldives Bahrain
Cirrhosis and other chronic liver diseases Qatar 0.162 0.026 0.03 United Arab Emirates Bahrain Maldives
Inflammatory bowel disease Oman 0.634 0.401 <0.001 Saudi Arabia Maldives Bahrain
Alzheimer’s disease and other dementias Japan -0.925 0.855 <0.001 Italy Portugal Greece
Parkinson’s disease Japan -0.936 0.877 <0.001 Italy Portugal Greece
Alcohol use disorders Oman 0.438 0.192 <0.001 Saudi Arabia Maldives Bahrain
Diabetes mellitus Sri Lanka -0.355 0.126 <0.001 Saint Vincent and the Grenadines Brazil Argentina
Chronic kidney disease Uruguay -0.411 0.169 <0.001 New Zealand Iceland Australia
Rheumatoid arthritis Puerto Rico -0.713 0.509 <0.001 Portugal Croatia Finland
Maternal disorders Central African Republic -0.909 0.827 <0.001 Niger Somalia Chad
Neonatal disorders Central African Republic -0.91 0.827 <0.001 Niger Somalia Chad
Self-harm United Arab Emirates 0.443 0.197 <0.001 Qatar Bahrain Maldives
Interpersonal violence Tonga -0.472 0.223 <0.001 Samoa Syria Comoros
HIV/AIDS Central African Republic -0.549 0.302 <0.001 Niger Somalia Chad
Tuberculosis Malta 0.759 0.576 <0.001 Poland Cyprus Slovakia
Dengue Jamaica -0.578 0.334 <0.001 Colombia Saint Lucia Brunei
Protein-energy malnutrition Malta 0.561 0.315 <0.001 Poland Cyprus Slovakia
3.2.3 Level 3 Disease Categories
Table 1, presents PoPStat related to selected Level 3 disease categories, along with their corresponding reference countries, countries adjacent to reference populations and other important statistical measures.
@@ 4.2 Performance of PoPStat with other indicators
Examining IHD as an illustrative example, the association of IHD mortality with PoPDivergence - represented by the PoPStat - was superior to the correlations of other demographic indicators included in this study. This pattern remained true for all the relevant causes and cause clusters across the three levels (Supplemental File 1). This indicates that other scalar variables such as median age and life expectancy may condense lesser information about the demography and its health compared to the PoPDivergence. Therefore, PopDivergance becomes a high-quality dimension reduction of the information-rich population pyramid. For example, in relation to Figure 2(a), PoPDivergence captures the shape differences of the two population pyramids that share a singular median age. However, it must be noted that the correlation between median age and disease mortality remained equally strong for a majority of causes (Supplemental File 1). So, future research can explore if the ability of these two indicators to associate with the underlying demographic structures of the demographic transition remains true for areas other than epidemiology.

### v2

@@ 3 Results
(a) Non-Communicable Diseases
(b) Communicable, Maternal, Neonatal, and Nutritional Diseases
(c) Injuries
(d) Neurological Disorders
(e) Neglected Tropical Diseases
(f) Transport Injuries
(g) Cardiovascular Diseases
(h) Respiratory Infections and Tuberculosis
Figure 4: PoPDiveregence vs Disease Mortality for Selected High Impact Disease Categories
PoPStat values were calculated for all disease categories up to Level 3 are summarized in Table 1 and 2. Figure 4, shows variations of PoPDivergence and mortality rates for selected disease categories, where population pyramids of randomly selected countries are shown for each plot to illustrate the variation of PoPDivergence.
At Level 1, among the three aggregated disease categories, non communicable diseases (NCD) exhibited the highest association with overall mortality, with a PoPStat of (–0.84, Japan, p<0.001), explaining 81.4% of the mortality variation. As shown in Figure 4(a) countries with constrictive and more symmetrically shaped population pyramids such as Japan, Italy, Portugal, and Greece had higher mortality due to NCDs. In contrast, the aggregate of communicable, maternal, neonatal, and nutritional diseases (CMNN) showed a moderate association with a PoPStat of (0.50, Singapore, p<0.001), accounting for 25.4% of the variation as shown in Figure 4(b). Countries with more symmetrical and constrictive pyramids like Singapore, Mauritius, Chile, and Saint Lucia saw lower CMNN death rates. Injuries, as visualized in Figure 4(c), exhibited the weakest association, with a PoPStat of (0.29, Singapore, p<0.001), explaining only 8.4% of the variation in injury-related mortality. Countries like Singapore, Mauritius, Chile, and Saint Lucia with constrictive pyramids that were largely symmetrical had lower injury related mortality.
Within the NCD group at Level 2, neurological diseases recorded the highest PoPStat of (0.90, Zambia, p<0.001), explaining 81.4% of the mortality variation attributed to neurological causes as depicted in Figure 4(d). Countries like Burkina Faso, Cameroon, and Yemen, with expansive, predominantly symmetrical population pyramids, had a lower burden of deaths due to neurological diseases. Within the CMNN group at Level 2, under communicable diseases, neglected tropical diseases (NTD) had the highest PoPStat of (0.83, Malta, p<0.001), which explained 69.7% of the NTD-related mortality variation, as visualized in Figure 4(e). Countries with symmetrical, constrictive population pyramids such as Malta, Poland, Cyprus, and Slovakia experienced lower mortality from NTDs. Under maternal, neonatal, and nutritional diseases, maternal and neonatal conditions exhibited the highest PoPStat(0.91, Central African Republic, p<0.001), explaining 83.6% of the variance in maternal and neonatal mortality. Countries such as Niger, Somalia, and Chad which shared expansive population pyramids saw higher death rates. Within the injuries category, transport injuries (TI), as shown in Figure 4(f) indicated a moderate association with a PoPStat of (–0.62, Burundi, p<0.001), explaining 38.5% of the mortality variation. Countries like Mali, Burkina Faso, and Mozambique showed similar symmetrically expansive pyramids and higher mortality.
At Level 2, focusing on the most severe subcategories, cardiovascular diseases (CVD) as the leading contributor among NCDs, exhibits a strong association of (–0.73, Japan, p<0.001), explaining 53.4% of the mortality variation. Figure 4(g), illustrates how countries with symmetrical and constrictive population pyramids identified before had more death rates due to cardiovascular diseases. Conversely, within the CMNN category, respiratory infections and tuberculosis showed a weak association, with a PoPStat of (0.143, Qatar, p=0.055) that explained only 2% of the mortality variation. As seen in 4(h), asymmetrically constrictive pyramids like Qatar, UAE, Bahrain, and Maldives had a lower burden of deaths.
Diving further on the most severe diseases in the CVD subgroup at Level 3, ischemic heart disease (IHD) demonstrated a robust association with a PoPStat of (–0.75, Latvia, p<0.001), explaining 56.2% of IHD related mortality. We use IHD as an example to calculate the correlations of other well recognized demographic and socio economic indicators with mortality rates and to compare their performance with PoPStat, as seen in 5(a).
@@
Table 1: PoPStat and Reference Populations for Level 1 Diseases and Selected Level 2 Diseases
Disease Category PoPStat CI PoPStat Closest countries to the reference
r Reference country p r2 First Second Third
Level 1 Disease Categories
Non-communicable diseases -0.846 Japan < 0.001 (-0.883, -0.799) 0.716 Italy Portugal Greece
@@
Communicable , maternal , neonatal, and nutritional diseases 0.504 Singapore < 0.001 (0.387, 0.605) 0.254 Mauritius Chile Saint Lucia
Injuries 0.291 Singapore < 0.001 (0.151, 0.419) 0.084 Mauritius Chile Saint Lucia
Level 2 Disease Categories
Non Communicable Diseases
Neurological disorders 0.902 Zambia < 0.001 (0.871, 0.926) 0.814 Burkina Faso Cameroon Yemen
Neoplasms -0.897 Japan < 0.001 (-0.922, -0.864) 0.804 Italy Portugal Greece
Cardiovascular diseases -0.731 Japan < 0.001 (-0.792, -0.655) 0.534 Italy Portugal Greece
Mental disorders -0.7 Japan < 0.001 (-0.767, -0.617) 0.49 Italy Portugal Greece
Musculoskeletal disorders -0.646 Japan < 0.001 (-0.724, -0.552) 0.418 Italy Portugal Greece
Chronic respiratory diseases -0.53 France < 0.001 (-0.627, -0.416) 0.281 Belgium United Kingdom Finland
Substance use disorders -0.533 Japan < 0.001 (-0.63, -0.419) 0.284 Italy Portugal Greece
Digestive diseases 0.499 Qatar < 0.001 (0.381, 0.601) 0.249 United Arab Emirates Bahrain Maldives
Other non-communicable diseases 0.487 Singapore < 0.001 (0.367, 0.591) 0.237 Mauritius Chile Saint Lucia
Diabetes and kidney diseases -0.408 Argentina < 0.001 (-0.523, -0.278) 0.166 Sri Lanka Costa Rica Saint Vincent and the Grenadines
@@
Neglected tropical diseases and malaria 0.835 Malta < 0.001 (0.785, 0.874) 0.697 Poland Cyprus Slovakia
Other infectious diseases 0.827 Malta < 0.001 (0.775, 0.869) 0.685 Poland Cyprus Slovakia
Enteric infections 0.735 Malta < 0.001 (0.66, 0.796) 0.54 Poland Cyprus Slovakia
HIV/AIDS and sexually transmitted infections -0.644 Central African Republic < 0.001 (-0.722, -0.55) 0.415 Niger Somalia Chad
Respiratory infections and tuberculosis 0.143 Qatar 0.055 (-0.003, 0.283) 0.02 United Arab Emirates Bahrain Maldives
Maternal , Neonatal, and Nutritional Diseases
Maternal and neonatal disorders -0.915 Central African Republic < 0.001 (-0.936, -0.887) 0.836 Niger Somalia Chad
Nutritional deficiencies 0.569 Cyprus < 0.001 (0.46, 0.66) 0.324 Luxembourg Chile Iceland
@@
Transport injuries -0.621 Burundi < 0.001 (-0.703, -0.522) 0.385 Mali Burkina Faso Mozambique
@@
Disease Category PoPStat CI PoPStat Closest countries to the reference
r Reference country p r2 First Second Third
Level 3 Disease Categories
Ischemic heart disease -0.75 Latvia < 0.001 (-0.807, -0.678) 0.562 Estonia Lithuania Slovenia
Stroke -0.497 France < 0.001 (-0.599, -0.379) 0.247 Belgium United Kingdom Finland
Pulmonary Arterial Hypertension 0.375 Oman < 0.001 (0.242, 0.494) 0.14 Saudi Arabia Maldives Bahrain
Chronic obstructive pulmonary disease -0.621 France < 0.001 (-0.704, -0.523) 0.386 Belgium United Kingdom Finland
Asthma -0.594 Central African Republic < 0.001 (-0.681, -0.49) 0.352 Niger Somalia Chad
Breast cancer -0.815 Japan < 0.001 (-0.859, -0.759) 0.664 Italy Portugal Greece
Colon and rectum cancer -0.922 Japan < 0.001 (-0.941, -0.897) 0.85 Italy Portugal Greece
Cervical cancer -0.257 Tonga < 0.001 (-0.388, -0.116) 0.066 Samoa Syria Comoros
Prostate cancer -0.716 Japan < 0.001 (-0.781, -0.637) 0.513 Italy Portugal Greece
Liver cancer 0.499 Oman < 0.001 (0.381, 0.601) 0.249 Saudi Arabia Maldives Bahrain
Cirrhosis and other chronic liver diseases 0.162 Qatar 0.03 (0.016, 0.3) 0.026 United Arab Emirates Bahrain Maldives
Inflammatory bowel disease 0.634 Oman < 0.001 (0.537, 0.714) 0.401 Saudi Arabia Maldives Bahrain
Alzheimer’s disease and other dementias -0.925 Japan < 0.001 (-0.943, -0.9) 0.855 Italy Portugal Greece
Parkinson’s disease -0.936 Japan < 0.001 (-0.952, -0.916) 0.877 Italy Portugal Greece
Alcohol use disorders 0.438 Oman < 0.001 (0.312, 0.549) 0.192 Saudi Arabia Maldives Bahrain
Diabetes mellitus -0.355 Sri Lanka < 0.001 (-0.476, -0.22) 0.126 Saint Vincent and the Grenadines Brazil Argentina
Chronic kidney disease -0.411 Uruguay < 0.001 (-0.525, -0.282) 0.169 New Zealand Iceland Australia
Rheumatoid arthritis -0.713 Puerto Rico < 0.001 (-0.778, -0.633) 0.509 Portugal Croatia Finland
Maternal disorders -0.909 Central African Republic < 0.001 (-0.932, -0.88) 0.827 Niger Somalia Chad
Neonatal disorders -0.91 Central African Republic < 0.001 (-0.932, -0.881) 0.827 Niger Somalia Chad
Self-harm 0.443 United Arab Emirates < 0.001 (0.318, 0.553) 0.197 Qatar Bahrain Maldives
Interpersonal violence -0.472 Tonga < 0.001 (-0.578, -0.351) 0.223 Samoa Syria Comoros
HIV/AIDS -0.549 Central African Republic < 0.001 (-0.644, -0.439) 0.302 Niger Somalia chad
Tuberculosis 0.759 Malta < 0.001 (0.689, 0.815) 0.576 Poland Cyprus Slovakia
Dengue -0.578 Jamaica < 0.001 (-0.687, -0.444) 0.334 Colombia Saint Lucia Brunei
Protein-energy malnutrition 0.561 Malta < 0.001 (0.452, 0.654) 0.315 Poland Cyprus Slovakia
@@ 4.2 Performance of PoPStat with other indicators
Examining IHD as an illustrative example, the association of IHD mortality with PoPDivergence - represented by the PoPStat - was superior to the correlations of other demographic indicators included in this study. This pattern remained true for all the relevant causes and cause clusters across the three levels (Supplemental File 1). This indicates that other scalar variables such as median age and life expectancy may condense lesser information about the demography and its health compared to the PoPDivergence. Therefore, PopDivergance becomes a high-quality dimension reduction of the information-rich population pyramid. For example, in relation to Figure 1(b), PoPDivergence captures the shape differences of the two population pyramids that share a singular median age. However, it must be noted that the correlation between median age and disease mortality remained equally strong for a majority of causes (Supplemental File 1). So, future research can explore if the ability of these two indicators to associate with the underlying demographic structures of the demographic transition remains true for areas other than epidemiology.